OpenStudy (anonymous):

REALLY EASY PLZ HELP!!! MEDAL AND FAN PPLZ!!! I need one linear, exponential, and quadratic equation that all intersect in ++ on the graph and where both x and y are less than 10

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@Phebe @uri @equineheart

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@amistre64 @TheSmartOne @shifuyanli @AdamK

OpenStudy (anonymous):

grrrrrrrrr

OpenStudy (anonymous):

@GeniousCreation @danish071996 @DarkBlueChocobo

OpenStudy (anonymous):

@nincompoop @pooja195

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

Noooooo

OpenStudy (anonymous):

@triciaal

OpenStudy (campbell_st):

start with something easy... y = x^2 and y = x use this site to check they intersect in the 1st quadrant... https://www.desmos.com/ then for the cubic just play with it

OpenStudy (anonymous):

It's on paper so there's no graph that I can show you

OpenStudy (campbell_st):

and a 3rd curve might be y = x^3 - x or y = x^3 - x^2

OpenStudy (campbell_st):

use the graphing site to see if all the equations intersect in the 1st quadrant

OpenStudy (anonymous):

I only got the last one

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (campbell_st):

all 3 curves I posted go through the 1st quadrant... so graph them and see what happens

OpenStudy (anonymous):

No I only got the last curve

OpenStudy (anonymous):

plus I need one quadratic, one exponential, and one linear so they can't all be curves

OpenStudy (campbell_st):

use a basic parabola... y = x^2 and for the linear equation y = x

OpenStudy (anonymous):

okay... That works I guess so thanks!